# and Tight DP Auditing

### v1

@@ 3.3 Desiderata: The Gold Standard of DP Auditing
| Ding et al. [34] | Simple | -DP | – | – | Output-set | ✗ | ✓ | ✓ |
| Bischel et al. [18] | Simple | -DP | – | – | Output-set | ✗ | ✓ | ✓ |
| Jagielski et al. [54] | DP-SGD | ()-DP | Poisson | MI | FPR/FNR | ✗ | ✓ | ✗ |
| Nasr et al. [83] | DP-SGD | ()-DP | Poisson | MI | FPR/FNR | ✗ | ✓ | ✗ |
| Malek et al. [72] | ALIBI | Label DP | – | AI | Accuracy | ✓ | ✓ | ✗ |
| Houssiau et al. [52] | Synthetic Data | ()-DP | – | AI | FPR/FNR | ✗ | ✓ | ✗ |
| Askin et al. [12] | Simple | -DP | – | – | Distance Est. | ✗ | ✓ | ✓ |
| Lokna et al. [66] | Simple | ()-DP | – | DPD | Output-set | ✗ | ✓ | ✓ |
| Nasr et al. [82] | DP-SGD | -DP | Poisson | MI | FPR/FNR | ✗ | ✓ | ✗ |
| Maddock et al. [70] | DP-FedSGD | User-level DP | Poisson | MI | FPR/FNR | ✓ | ✗ | ✗ |
| Pillutla et al. [91] | DP-SGD | ()-DP | Poisson | MI | Custom | ✗ | ✓ | ✗ |
| Steinke et al. [100] | DP-SGD | ()-DP | Poisson | MI | Accuracy | ✓ | ✓ | ✗ |
| Galen et al. [4] | DP-FedAvg | User-level DP | Shuffle | MI | Variance Est. | ✓ | ✗ | ✗ |
| Chadha et al. [21] | PATE | -RDP | – | MI | Custom | ✗ | ✗ | ✗ |
| Annamalai et al. [8] | Synthetic Data | ()-DP | Poisson | MI | FPR/FNR | ✗ | ✓ | ✗ |
| Feng et al. [45] | DP-SGD | ()-DP | Poisson | – | Output-set | ✗ | ✓ | ✓ |
| Annamalai et al. [7] | DP-SGD | -GDP | – | MI | FPR/FNR | ✗ | ✓ | ✗ |
| Mahloujifar et al. [71] | DP-SGD | -DP | Poisson | Recon | Accuracy | ✓ | ✗ | ✗ |
@@
| Annamalai et al. [6] | DP-SGD (Shuffle) | ()-DP | Shuffle | MI | FPR/FNR | ✗ | ✗ | ✗ |
| Cebere et al. [20] | DP-SGD | -GDP | Poisson | MI | FPR/FNR | ✗ | ✓ | ✗ |
| Xiang et al. [111] | DP-SGD | -DP | Poisson | MI | Accuracy | ✓ | ✗ | ✗ |
@@ 6.2 Sub-sampling Schemes
In this context, the main challenge is having tight DP guarantees of mechanisms employing sub-sampling schemes, i.e., knowing that the actual privacy leakage of mechanisms matches the theoretical privacy leakage expected from the DP guarantee. Next, we focus on the Gaussian Mechanism since sub-sampling schemes are most commonly used with that, e.g., in DP-SGD [1]. In Table 2, we summarize the sub-sampling schemes and adjacency notions for which tight DP guarantees are currently known. While tight guarantees are known for Poisson sub-sampling, Sampling WR, and Balls-in-Bins sampling for the Add/Remove, Edit, and Zero-out adjacencies, respectively, for the vast majority of sub-sampling scheme and adjacency pairs, tight guarantees remain unknown.
@@ Appendix B Prior Work on DP auditing
| Ding et al. [34] | Report Noisy Max Noisy Hist SVT | -DP | – | – | Output-set | ✗ | ✓ | ✓ |
| Bischel et al. [18] | Report Noisy Max NoisySum AboveThreshold | -DP | – | – | Output-set | ✗ | ✓ | ✓ |
| Liu et al. [65] | Report Noisy Max Noisy Hist SVT Truncated Geometric | ()-DP | – | – | Output-set | ✗ | ✓ | ✓ |
| Jagielski et al. [54] | DP-SGD | ()-DP | Poisson | MI | FPR/FNR | ✗ | ✓ | ✗ |
@@
| Bischel et al. [19] | Report Noisy Max Noisy Hist SVT RAPPOR Prefix Sum Truncated Geometric | ()-DP | – | DPD | Output-set | ✗ | ✓ | ✓ |
| Bernau et al. [17] | DP-SGD | ()-DP | Poisson | DPD MI | Output-set | ✗ | ✗ | ✓ |
| Wang et al. [107] | Laplace | -DP | – | DPD | Output-set | ✗ | ✓ | ✓ |
| Nasr et al. [83] | DP-SGD | ()-DP | Poisson | MI | FPR/FNR | ✗ | ✓ | ✗ |
| Malek et al. [72] | PATE-FM ALIBI | Label DP | – | AI | Accuracy | ✓ | ✓ | ✗ |
| Lu et al. [67] | Naive Bayes Random Forest DP-SGD | ()-DP | Poisson | MI | FPR/FNR | ✗ | ✓ | ✗ |
| Houssiau et al. [52] | CTGAN MST PrivBayes | ()-DP | – | AI | FPR/FNR | ✗ | ✓ | ✗ |
| Niu et al. [86] | Report Noisy Max NoisyHist SVT Laplace RAPPOR | ()-DP | – | – | Output-set | ✗ | ✓ | ✓ |
| Tramèr et al. [104] | Backpropagation Clipping | ()-DP | Poisson | MI | FPR/FNR | ✗ | ✓ | ✗ |
| Askin et al. [12] | Report Noisy Max SVT Laplace Exponential Mechanism | -DP | – | – | Distance Est. | ✗ | ✓ | ✓ |
| Lokna et al. [66] | MST Gaussian Discrete Gaussian Laplace | ()-DP | – | DPD | Output-set | ✗ | ✓ | ✓ |
| Zanella-Béguelin et al. [115] | DP-SGD | ()-DP | Poisson | MI | FPR/FNR | ✗ | ✓ | ✗ |
| Nasr et al. [82] | DP-SGD | -DP | Poisson | MI | FPR/FNR | ✗ | ✓ | ✗ |
| Gorla et al. [50] | Truncated Laplace | -RDP | – | – | Histogram Est. | ✗ | ✓ | ✓ |
| Maddock et al. [70] | DP-FedSGD | User-level DP | Poisson | MI | FPR/FNR | ✓ | ✗ | ✗ |
| Pillutla et al. [91] | DP-SGD | ()-DP | Poisson | MI | Custom | ✗ | ✓ | ✗ |
| Steinke et al. [100] | DP-SGD | ()-DP | Poisson | MI | Accuracy | ✓ | ✓ | ✗ |
| Galen et al. [4] | DP-FedAvg | User-level DP | Shuffle | MI | Variance Est. | ✓ | ✗ | ✗ |
| Matsumoto et al. [73] | LDP-SGD | -DP | Shuffle | MI | FPR/FNR | ✗ | ✗ | ✗ |
| Chadha et al. [21] | PATE CaPC PromptPATE Private k-NN | -RDP | – | MI | Custom | ✗ | ✗ | ✗ |
@@
| Zhang et al. [118] | Report Noisy Max NoisyHist SVT Laplace RAPPOR Prefix Sum Truncated Geometric | -DP | – | DPD | Output-set | ✗ | ✓ | ✓ |
| Chida et al. [24] | PrivBayes DPCopula Randomized Response | MI | – | MI | FPR/FNR | ✗ | ✓ | ✗ |
| Lu et al. [68] | Report Noisy Max NoisyHist SVT | -RDP | – | DPD | Output-set | ✗ | ✓ | ✓ |
| Xiang et al. [110] | DP-GNN | -DP | Poisson | MI | FPRFNR | ✓ | ✗ | ✓ |
| Annamalai et al. [8] | PrivBayes CTGAN DP-WGAN | ()-DP | Poisson | MI | FPR/FNR | ✗ | ✓ | ✗ |
| Feng et al. [45] | DP-SGD | ()-DP | Poisson | – | Output-set | ✗ | ✓ | ✓ |
| Annamalai et al. [5] | DP-SGD | -DP | Poisson | MI | FPRFNR | ✗ | ✗ | ✓ |
| Annamalai et al. [7] | DP-SGD | -GDP | – | MI | FPR/FNR | ✗ | ✓ | ✗ |
| Yoon et al. [114] | DP-SGD | -GDP | – | MI | FPR/FNR | ✗ | ✓ | ✗ |
| Kong et al. [62] | Laplace Gaussian SVT DP-SGD | ()-DP | – | – | Divergence Est. | ✗ | ✗ | ✓ |
| Debenedetti et al. [31] | DP-SGD | -DP | Poisson | MI | FPR/FNR | ✗ | ✗ | ✗ |
@@
| Arcolezi et al. [10] | Generalized Randomized Response Subset Selection Local Hashing Unary Encoding Histogram Encoding | ()-DP | – | DPD | Output-set | ✗ | ✓ | ✓ |
| Mahloujifar et al. [71] | DP-SGD | -DP | Poisson | Recon | Accuracy | ✓ | ✗ | ✗ |
@@
| Annamalai et al. [6] | DP-SGD (Shuffle) | ()-DP | Shuffle | MI | FPR/FNR | ✗ | ✗ | ✗ |
| Huang et al. [53] | Laplace Gaussian DP-SGD | ()-DP | Poisson | – | Variance Est. | ✗ | ✗ | ✓ |
| Nasr et al. [84] | DP-SGD | -DP | Poisson | MI | FPR/FNR | ✗ | ✓ | ✗ |
| Cebere et al. [20] | DP-SGD | -GDP | Poisson | MI | FPR/FNR | ✗ | ✓ | ✗ |
| Ganev et al. [47] | PATE-GAN | ()-DP | – | MI | FPR/FNR ✗ | ✓ | ✗ | |
| Askin et al. [11] | Laplace Gaussian Sub-Sampling DP-SGD | -DP | Poisson | DPD | Output-set | ✗ | ✓ | ✗ |
| Panda et al. [88] | DP-SGD | ()-DP | Poisson | MI | Accuracy | ✓ | ✓ | ✗ |
| Xiang et al. [111] | DP-SGD Gaussian | -DP -GDP | Poisson – | MI | Accuracy | ✓ ✓ | ✗ ✓ | ✗ ✓ |

### v2

@@ 3.3 Desiderata: The Gold Standard of DP Auditing
| Ding et al. (2018) [33] | Simple | -DP | – | – | Output-set | ✗ | ✓ | ✓ |
| Bischel et al. (2018) [17] | Simple | -DP | – | – | Output-set | ✗ | ✓ | ✓ |
| Jagielski et al. (2018) [53] | DP-SGD | ()-DP | Poisson | MI | FPR/FNR | ✗ | ✓ | ✗ |
| Nasr et al. (2021) [84] | DP-SGD | ()-DP | Poisson | MI | FPR/FNR | ✗ | ✓ | ✗ |
| Malek et al. (2021) [73] | ALIBI | Label DP | – | AI | Accuracy | ✓ | ✓ | ✗ |
| Houssiau et al. (2022) [51] | Synthetic Data | ()-DP | – | AI | FPR/FNR | ✗ | ✓ | ✗ |
| Askin et al. (2022) [12] | Simple | -DP | – | – | Distance Est. | ✗ | ✓ | ✓ |
| Lokna et al. (2023) [67] | Simple | ()-DP | – | DPD | Output-set | ✗ | ✓ | ✓ |
| Nasr et al. (2023) [83] | DP-SGD | -DP | Poisson | MI | FPR/FNR | ✗ | ✓ | ✗ |
| Maddock et al. (2023) [71] | DP-FedSGD | User-level DP | Poisson | MI | FPR/FNR | ✓ | ✗ | ✗ |
| Pillutla et al. (2023) [92] | DP-SGD | ()-DP | Poisson | MI | Custom | ✗ | ✓ | ✗ |
| Steinke et al. (2023) [101] | DP-SGD | ()-DP | Poisson | MI | Accuracy | ✓ | ✓ | ✗ |
| Galen et al. (2024) [4] | DP-FedAvg | User-level DP | Shuffle | MI | Variance Est. | ✓ | ✗ | ✗ |
| Chadha et al. (2024) [20] | PATE | -RDP | – | MI | Custom | ✗ | ✗ | ✗ |
| Annamalai et al. (2024) [8] | Synthetic Data | ()-DP | Poisson | MI | FPR/FNR | ✗ | ✓ | ✗ |
| Feng et al. (2024) [44] | DP-SGD | ()-DP | Poisson | – | Output-set | ✗ | ✓ | ✓ |
| Annamalai et al. (2024) [7] | DP-SGD | -GDP | – | MI | FPR/FNR | ✗ | ✓ | ✗ |
| Mahloujifar et al. (2024) [72] | DP-SGD | -DP | Poisson | Recon | Accuracy | ✓ | ✗ | ✗ |
@@
| Annamalai et al. (2024) [6] | DP-SGD (Shuffle) | ()-DP | Shuffle | MI | FPR/FNR | ✗ | ✗ | ✗ |
| Cebere et al. (2025) [19] | DP-SGD | -GDP | Poisson | MI | FPR/FNR | ✗ | ✓ | ✗ |
| Xiang et al. (2025) [112] | DP-SGD | -DP | Poisson | MI | Accuracy | ✓ | ✗ | ✗ |
@@ 6.2 Sub-sampling Schemes
In this context, the main challenge is having tight DP guarantees of mechanisms employing sub-sampling schemes, i.e., knowing that the actual privacy leakage of mechanisms matches the theoretical privacy leakage expected from the DP guarantee. Next, we focus on the Gaussian Mechanism since sub-sampling schemes are most commonly used with that, e.g., in DP-SGD [1]. In Table 2, we summarize the sub-sampling schemes and adjacency notions for which tight DP guarantees are currently known. While tight guarantees are known for Poisson sub-sampling for any adjacency and Sampling WOR, Sampling WR, and Balls-in-Bins sampling for the Add/Remove, Edit, and Zero-out adjacencies, respectively, for the vast majority of sub-sampling scheme and adjacency pairs, tight guarantees remain unknown.
@@ Appendix B Prior Work on DP auditing
| Ding et al. (2018) [33] | Report Noisy Max Noisy Hist SVT | -DP | – | – | Output-set | ✗ | ✓ | ✓ |
| Bischel et al. (2018) [17] | Report Noisy Max NoisySum AboveThreshold | -DP | – | – | Output-set | ✗ | ✓ | ✓ |
| Liu et al. (2019) [66] | Report Noisy Max Noisy Hist SVT Truncated Geometric | ()-DP | – | – | Output-set | ✗ | ✓ | ✓ |
| Jagielski et al. (2020) [53] | DP-SGD | ()-DP | Poisson | MI | FPR/FNR | ✗ | ✓ | ✗ |
@@
| Bischel et al. (2021) [18] | Report Noisy Max Noisy Hist SVT RAPPOR Prefix Sum Truncated Geometric | ()-DP | – | DPD | Output-set | ✗ | ✓ | ✓ |
| Bernau et al. (2021) [16] | DP-SGD | ()-DP | Poisson | DPD MI | Output-set | ✗ | ✗ | ✓ |
| Wang et al. (2021) [108] | Laplace | -DP | – | DPD | Output-set | ✗ | ✓ | ✓ |
| Nasr et al. (2021) [84] | DP-SGD | ()-DP | Poisson | MI | FPR/FNR | ✗ | ✓ | ✗ |
| Malek et al. (2021) [73] | PATE-FM ALIBI | Label DP | – | AI | Accuracy | ✓ | ✓ | ✗ |
| Lu et al. (2022) [68] | Naive Bayes Random Forest DP-SGD | ()-DP | Poisson | MI | FPR/FNR | ✗ | ✓ | ✗ |
| Houssiau et al. (2022) [51] | CTGAN MST PrivBayes | ()-DP | – | AI | FPR/FNR | ✗ | ✓ | ✗ |
| Niu et al. (2022) [87] | Report Noisy Max NoisyHist SVT Laplace RAPPOR | ()-DP | – | – | Output-set | ✗ | ✓ | ✓ |
| Tramèr et al. (2022) [105] | Backpropagation Clipping | ()-DP | Poisson | MI | FPR/FNR | ✗ | ✓ | ✗ |
| Askin et al. (2022) [12] | Report Noisy Max SVT Laplace Exponential Mechanism | -DP | – | – | Distance Est. | ✗ | ✓ | ✓ |
| Lokna et al. (2023) [67] | MST Gaussian Discrete Gaussian Laplace | ()-DP | – | DPD | Output-set | ✗ | ✓ | ✓ |
| Zanella-Béguelin et al. (2023) [116] | DP-SGD | ()-DP | Poisson | MI | FPR/FNR | ✗ | ✓ | ✗ |
| Nasr et al. (2023) [83] | DP-SGD | -DP | Poisson | MI | FPR/FNR | ✗ | ✓ | ✗ |
| Gorla et al. (2023) [49] | Truncated Laplace | -RDP | – | – | Histogram Est. | ✗ | ✓ | ✓ |
| Maddock et al. (2023) [71] | DP-FedSGD | User-level DP | Poisson | MI | FPR/FNR | ✓ | ✗ | ✗ |
| Pillutla et al. (2023) [92] | DP-SGD | ()-DP | Poisson | MI | Custom | ✗ | ✓ | ✗ |
| Steinke et al. (2023) [101] | DP-SGD | ()-DP | Poisson | MI | Accuracy | ✓ | ✓ | ✗ |
| Galen et al. (2024) [4] | DP-FedAvg | User-level DP | Shuffle | MI | Variance Est. | ✓ | ✗ | ✗ |
| Matsumoto et al. (2024) [74] | LDP-SGD | -DP | Shuffle | MI | FPR/FNR | ✗ | ✗ | ✗ |
| Foundations | Operational Details | Progress | ||||||
| Chadha et al. (2024) [20] | PATE CaPC PromptPATE Private k-NN | -RDP | – | MI | Custom | ✗ | ✗ | ✗ |
@@
| Zhang et al. (2024) [119] | Report Noisy Max NoisyHist SVT Laplace RAPPOR Prefix Sum Truncated Geometric | -DP | – | DPD | Output-set | ✗ | ✓ | ✓ |
| Chida et al. (2024) [23] | PrivBayes DPCopula Randomized Response | MI | – | MI | FPR/FNR | ✗ | ✓ | ✗ |
| Lu et al. (2024) [69] | Report Noisy Max NoisyHist SVT | -RDP | – | DPD | Output-set | ✗ | ✓ | ✓ |
| Xiang et al. (2024) [111] | DP-GNN | -DP | Poisson | MI | FPRFNR | ✓ | ✗ | ✓ |
| Annamalai et al. (2024) [8] | PrivBayes CTGAN DP-WGAN | ()-DP | Poisson | MI | FPR/FNR | ✗ | ✓ | ✗ |
| Feng et al. (2024) [44] | DP-SGD | ()-DP | Poisson | – | Output-set | ✗ | ✓ | ✓ |
| Annamalai et al. (2024) [5] | DP-SGD | -DP | Poisson | MI | FPRFNR | ✗ | ✗ | ✓ |
| Annamalai et al. (2024) [7] | DP-SGD | -GDP | – | MI | FPR/FNR | ✗ | ✓ | ✗ |
| Yoon et al. (2024) [115] | DP-SGD | -GDP | – | MI | FPR/FNR | ✗ | ✓ | ✗ |
| Kong et al. (2024) [61] | Laplace Gaussian SVT DP-SGD | ()-DP | – | – | Divergence Est. | ✗ | ✗ | ✓ |
| Debenedetti et al. (2024) [30] | DP-SGD | -DP | Poisson | MI | FPR/FNR | ✗ | ✗ | ✗ |
@@
| Arcolezi et al. (2024) [10] | Generalized Randomized Response Subset Selection Local Hashing Unary Encoding Histogram Encoding | ()-DP | – | DPD | Output-set | ✗ | ✓ | ✓ |
| Mahloujifar et al. (2024) [72] | DP-SGD | -DP | Poisson | Recon | Accuracy | ✓ | ✗ | ✗ |
@@
| Annamalai et al. (2024) [6] | DP-SGD (Shuffle) | ()-DP | Shuffle | MI | FPR/FNR | ✗ | ✗ | ✗ |
| Huang et al. (2025) [52] | Laplace Gaussian DP-SGD | ()-DP | Poisson | – | Variance Est. | ✗ | ✗ | ✓ |
| Nasr et al. (2025) [85] | DP-SGD | -DP | Poisson | MI | FPR/FNR | ✗ | ✓ | ✗ |
| Cebere et al. (2025) [19] | DP-SGD | -GDP | Poisson | MI | FPR/FNR | ✗ | ✓ | ✗ |
| Ganev et al. (2025) [46] | PATE-GAN | ()-DP | – | MI | FPR/FNR | ✗ | ✓ | ✗ |
| Askin et al. (2025) [11] | Laplace Gaussian Sub-Sampling DP-SGD | -DP | Poisson | DPD | Output-set | ✗ | ✓ | ✗ |
| Panda et al. (2025) [89] | DP-SGD | ()-DP | Poisson | MI | Accuracy | ✓ | ✓ | ✗ |
| Xiang et al. (2025) [112] | DP-SGD Gaussian | -DP -GDP | Poisson – | MI | Accuracy | ✓ ✓ | ✗ ✓ | ✗ ✓ |